# Characterization of the directed landscape from the KPZ fixed point

### v2

@@ 1.3 Structure of the paper and proof overview
After a short preliminary section, we prove Theorem 1.1 in Section 3, up to a key technical result that we leave for Section 4. These are the most important parts of the paper. We then prove Theorem 1.3 from Theorem 1.1 in Section 5. This section also contains a technical extension of Theorem 1.3 that we will use to study AEPs, where the known convergence to the KPZ fixed point is slightly more delicate. Finally, Sections 6, 7 and 8 prove the applications in Theorem 1.6. The proofs in the final sections are mostly straightforward, though there are some technical lemmas required to handle monotonicity of the more delicate models (in particular, general AEPs).
@@
This is fairly quick, and we refer the reader to 3.1 for details. Given this structure, to show that =$\mathcal{M}\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}$, it suffices to prove that $\mathcal{M}_{s,t}$ and $\mathcal{L}_{s,t}$ have the same finite dimensional distributions for all fixed $s<t$. We will outline how we check that for a set $F=\{(x_{1},y_{1}),(x_{2},y_{2})\}$, the two-point distribution $\mathcal{M}_{0,1}|_{F}$ matches that of $\mathcal{L}$. The general case uses the same idea. For $s\in[0,1]$ define $\mathcal{M}^{s}=\mathcal{L}_{0,s}\diamond\mathcal{M}_{s,1},$ so that $\mathcal{M}^{0}=\mathcal{M}_{0,1}$ and $\mathcal{M}^{1}=\mathcal{L}_{0,1}$. We will aim to show that with respect to the $L^{1}$-Wasserstein metric $d_{W}$, for $s<t=s+r$,
@@
This type of compatibility can be viewed as roughly corresponding to the event where in a candidate landscape $\mathcal{M}^{t}$, if $\pi_{i},i=1,2$ are the geodesics from $(x_{i},0)$ and $(y_{i},1)$, then either $\pi_{1}(t),\pi_{2}(t)$ are far away from each other ($A_{1},A_{2}$ are separated), or else $\pi_{1}(t)=\pi_{2}(t)$ and the geodesics coalesce together well before time $s$ ($A_{1},A_{2}$ overlap and the functions $\mathcal{L}_{0,s}(x_{1},\cdot),\mathcal{L}_{0,s}(x_{2},\cdot)$ differ by a constant). In other words, the coalescence time for the geodesics $\pi_{1},\pi_{2}$ occurs far away from the interval $[s,t]$. If $\mathcal{M}^{s},\mathcal{M}^{t}$ were true directed landscapes, we could prove that this compatibility event holds with probability $1-r^{1-o(1)}$. For pre-landscapes, we will show the weaker bound $1-r^{2/3+\delta}$, which still gives (10). In practice, we will also work with a stationary version of $\mathcal{M}^{s}|_{F},\mathcal{M}^{t}|_{F}$ in order to facilitate technical computations.
@@
Acknowledgments. The authors thank Jeremy Quastel for several useful discussions regarding convergence of general AEPs. The authors also thank Amol Aggarwal for pointing out the argument in [ACH24b, Appendix D.3], and the references [BGRS02, BGRS19], and Ivan Corwin, Milind Hegde, and Shalin Parekh for useful comments on an earlier version of the paper.
@@ 2.1 The directed landscape and the KPZ fixed point
Recall that $\operatorname{UC}$ is the set of all upper semicontinuous functions from $\mathbb{R}\to\bar{\mathbb{R}}$. Then a function $f$ is in $\operatorname{UC}$ if and only if its hypograph
@@
in the Hausdorff metric on closed subsets of ×[1,$\mathbb{R}\times[-1,1]$, where $E:\mathbb{R}\times\bar{\mathbb{R}}\to\mathbb{R}\times[-1,1]$ is the map $E(x,s)=(x,e^{-|x|}s/(1+|s|)),$ and the $d_{H}$ metric on the right-hand side above is the usual Hausdorff metric. It is straightforward to check that $\operatorname{UC}$ is a compact Polish space. Setting some notations, we write $\delta_{x}\in\operatorname{UC}$ for the narrow wedge at $x$, i.e., the function with $\delta_{x}(x)=0$ and $\delta_{x}(y)=-\infty$ for $y\neq x$. For $f\in\operatorname{UC}$ and $K\subset\mathbb{R}$ we also write $f|_{K}\in\operatorname{UC}$ for the function equal to $f$ on $K$, and $\infty$- ∞ on $K^{c}$.
@@
[…]
@@
As already alluded to, we can alternately describe the KPZ fixed point directly in terms of the directed landscape. Indeed, for $f\in\operatorname{UC}$, define $\mathfrak{h}_{0}f=f$, and $\mathfrak{h}_{t}f$ for $t>0$ through (4) (i.e., $\mathfrak{h}_{t}f:=f\diamond\mathcal{L}_{0,t}$);
[…]
@@ 2.2 Basic properties of the directed landscape
We say that $d$ is a directed metric of negative sign if $-d$ is a directed metric of positive sign. Many natural KPZ models are directed metrics but not metrics, and moreover, unlike the class of metrics, the class of directed metrics is preserved under the kinds of rescaling typical seen in the KPZ class. If we extend the directed landscape to equal $\infty$- ∞ on $\mathbb{R}^{4}\setminus\mathbb{R}^{4}_{\uparrow}$ then it becomes a directed metric (of negative sign) on $\mathbb{R}^{2}$.
@@ 2.3 The stationary horizon
Here $f[i\to x]$ is a last passage value from $(i,-\infty)$ to $(x,1)$ in the sense of [DV24], and $|\pi|_{f}$ can be thought of as the length of a path associated to $\pi$. Now let $a_{1}<\dots<a_{k}$ and let =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{k}$ be a family of independent two-sided Brownian motions where each $\mathcal{B}_{i}$ has slope $a_{i}$. Define =$\mathcal{R}=\{\mathcal{R}_{i}\}_{i=1}^{k}$ by
@@ 3.1 Continuity and metric composition
on $\mathbb{Q}^{2}$. Using the continuity of $\mathcal{M}$, and Corollary 3.6 which controls the location of the argmax in the metric composition, we can deduce that both sides of (20) are continuous, and are also continuous in $s,r$, and $t$. Thus we can extend (20) to $\mathbb{R}^{2}$, and to hold for any $s<r<t$ simultaneously.
@@ 3.2 The law of $\mathcal{M}_{0,1}$ and an exchange strategy
(,$(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{M}_{0,1}(y_{i})\}_{i=1}^{k})\stackrel{{\scriptstyle d}}{{=}}(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{L}_{0,1}(y_{i})\}_{i=1}^{k}).$ (21)
@@
By 2.5 and Corollary 3.6, with probability at least $1-C\exp(-cM)$, the bound in 2.5 holds for both $\mathcal{M},\mathcal{L}$ with some $R<M^{2/3}$. Also, with probability $>1-C\exp(-cM)$, we have $|\mathcal{R}_{i}(x)-\mathcal{R}_{i}(0)-a_{i}x|<|x|+M$ for all $x\in\mathbb{R},i=1,\dots,k$. Let $\mathcal{E}$ denote the intersection of the above two events. Then on $\mathcal{E}$, we have that
@@ 3.3 The coupling
We now define the coupling to be used in 3.9. Below we fix $s<t=s+r\in[\theta,1-\theta]$, and throughout assume $r=t-s$ is sufficiently small. We start with independent copies of $\mathcal{R}$, $\mathcal{L}_{0,s}$, and $\mathcal{M}_{t,1}$, and consider the following subsets of $\mathbb{R}$. For each $i\in\llbracket 1,k\rrbracket$, denote $H_{i}=\mathcal{R}_{i}\diamond\mathcal{L}_{0,s},$ and
@@
We let $A_{i}$ be the $\log^{10}(r)r^{2/3}$-neighborhood of $\tilde{A}_{i}$. For each $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$, we let $\tilde{D}_{i,j}$ be the set where $H_{i}-H_{j}$ is non-constant, i.e.,
$\tilde{D}_{i,j}=\{x\in\mathbb{R}:\text{ for all }\epsilon>0,\text{ there exists }y\in(x-\epsilon,x+\epsilon)\text{ with }H_{i}(x)-H_{j}(x)\neq H_{i}(y)-H_{j}(y)\}.$
@@
There is a constant $\delta>0$, such that for all sufficiently small $r>0$ and $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$, we have $\mathbb{P}[A_{i}\cap A_{j}\cap D_{i,j}=\emptyset]>1-r^{2/3+\delta}$.
@@
Let $\mathcal{E}_{*}$ be the event where $A_{i}\cap A_{j}\cap D_{i,j}=\emptyset$, for each $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$. Note that for each $i\in\llbracket 1,k\rrbracket$, (using the shape bounds in 2.5, Corollary 3.6 on $\mathcal{L},\mathcal{M}$) $A_{i}$ consists of finitely many open intervals $A_{i,1},\dots,A_{i,\ell(i)}$, each of length at least $2\log^{10}(r)r^{2/3}$.
@@
On $\mathcal{E}_{*}^{c}$, set $F=0$ on $\bigcup_{i=1}^{k}A_{i}$ and $\infty$- ∞ off this set. Next, for any interval $A_{i,j}$ and a function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ , we say that $f$ is suitable on $A_{i,j}$, if $f-H_{i}$ is constant on $I$. Sort the intervals $A_{i,j},i=1,\dots,k,j=1,\dots,\ell(i)$ into a list $I_{1},\ldots,I_{n}$ so that the $I_{j}$ are ordered by their left endpoint (ordering arbitrarily if left endpoints agree). We define $F$ on $I_{1},\dots,I_{n}$ inductively.
@@
It remains to check the last condition. By Theorem 2.9 the law of each $H_{i}$ is a Brownian motion with drift $a_{i}$. Then for any $M>1$, with probability $>1-C\exp(-cM)$ we have that $|H_{i}(x)-H_{i}(y)|<M\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for each $i\in\llbracket 1,k\rrbracket$ and $x,y\in[-M,M]$. Also, by 2.5/Corollary 3.6, with probability $>1-C\exp(-cM)$, we have that $\bigcup_{i=1}^{k}A_{i}\subset[-M,M]$. Thus the last condition holds from the construction of $F$. The measurability is immediate from construction. ∎
@@ 3.4 Removing the stationary initial condition
As $\epsilon\to 0$, the second bracketed term above remains bounded even after we take a supremum over $x$, since $\mathcal{R}_{i}$ is a Brownian motion with drift $i\leq k$; while by straightforward algebra the first term goes to $\infty$- ∞ uniformly over $x$ with $|x-x_{i}|>\epsilon$. This yields (28). For the bound (29), observe that for any $f:\mathbb{R}\to\mathbb{R}^{n}$ and $x<x_{i}<y$ we have
@@
Using these inequalities with $f=\mathcal{B}+\phi_{\epsilon}$ yields (29) since $\mathcal{B}$ is continuous and each $\phi_{\epsilon,j}$ is non-decreasing on $[x_{i}-\epsilon,x_{i}]$ and non-increasing on $[x_{i},x_{i}+\epsilon]$. ∎
@@
We will show that (for some $\delta>0$) these two probabilities are $<r^{4/3-\delta}$ and $<r^{3\delta}$, respectively.
@@
~$\tilde{\mathcal{B}}_{2}(x)=\mathcal{B}_{1}(x)+\sup_{z\leq x}\mathcal{B}_{2}(z)-\mathcal{B}_{1}(z).$
@@
On the other hand, again using Lemma 2.8 and the stationarity of $x\mapsto\mathcal{A}_{2}(x)+x^{2}$, we can deduce that $\mathbb{P}[\mathcal{E}_{1}^{\prime}\cap\mathcal{E}_{2}^{\prime}]<Cr^{4/3}\log^{44}(r)$, by replacing $\mathcal{M}_{1}-\mathcal{M}_{1}(z_{+})$ on $[z_{+}-1,z_{+}+1]$ with a Brownian motion $\mathcal{B}_{3}$ with $\mathcal{B}_{3}(z_{+})=0$ with drift in $[-C|\log(r)|,C|\log(r)|]$, and using Lemma 4.3 below (with $X=\mathcal{B}_{3}(z_{+}-\cdot),Y=\mathcal{B}_{1}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),Z=\mathcal{B}_{2}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),\epsilon=2\log^{10}(r)r^{1/3}$ and $D=C|\log(r)|$).
@@
Take any $D>1$ and $10<\epsilon<$ 1. Let $X,Y,Z$ be three independent Brownian motions, each with drift in $[-D,D]$. Then we have
@@
since ~$\tilde{\mathcal{B}}_{2}\geq\mathcal{B}_{2}$. Now, define the event $\mathcal{E}_{3}^{\prime}$ by
@@
We again use Lemma 2.8 and the stationarity of Airy2 plus a parabola, to replace $\mathcal{M}_{2}^{\prime}$ with a Brownian motion with drift in $[-C|\log(r)|,C|\log(r)|]$, conditioned on an event of probability at least $cr^{10}$. Then by Lemma 4.5 below (with $D=C|\log(r)|,Y=\mathcal{M}_{2}^{\prime}(\cdot+z_{+})-\mathcal{M}_{2}^{\prime}(z_{+}),X=\mathcal{B}_{2}(\cdot+z_{+})-\mathcal{B}_{2}(z_{+})$, and $\epsilon=2\log^{10}(r)r^{1/3}$) the above probability is at most $Cr^{\delta}$ for some absolute constant $\delta>0$. Putting everything together gives that
@@
There exists $\delta>0$ such that the following holds. Take any $D>1$ and $10<\epsilon<$ 1, with $D<\epsilon^{-1/4}$. Let $X$ and $Y$ be two independent Brownian motions, each with drift in $[-D,D]$. Let $\mathcal{E}$ be an event measurable with respect to $Y$, such that $\mathbb{P}[\mathcal{E}]>\epsilon^{100}$. Then
@@
For each $i\in\llbracket 1,2m\rrbracket$, we let $W^{\prime}_{i}$ be a Brownian motion with drift $-2^{-(i+1)/2}D$, conditional on $W^{\prime}_{i}\leq 0$ in $[0,1]$. Fix $(i_{1},\cdots,i_{m})\in\Phi$. By the claim above, for each $j\in\llbracket 1,m\rrbracket$, $W^{\prime}_{i_{j}}$ is stochastically dominated by the process $x\mapsto 2^{(i_{j}+1)/2}X(2^{-i_{j}-1}x)$ conditional on $\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}^{\prime}$. Therefore
@@
for any $I$ of length $r^{2/3}$, $I\subset[-|\log(r)|,|\log(r)|]$. Taking a union over at most $3r^{-2/3}\log(r)$ covering $[-|\log(r)|,|\log(r)|]$ and using Lemma 4.1 yields the result. ∎

### v3

@@ 1.3 Structure of the paper and proof overview
After a short preliminary section, we prove Theorem 1.1 in Section 3, up to a key technical result that we leave for Section 4. These are the most important parts of the paper. We then prove Theorem 1.3 from Theorem 1.1 in Section 5. This section also contains a technical extension of Theorem 1.3 that we will use to study general exclusion processes in Theorem 7.7, where random initial data is required to reach the KPZ fixed point. Finally, Sections 6 and 7 prove the applications in Theorem 1.6. The proofs in the final sections are mostly straightforward, though there are some technical lemmas required to handle monotonicity in general exclusion processes.
@@
This is fairly quick, and we refer the reader to 3.1 for details. Given this structure, to show that $\mathcal{M}\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}$, it suffices to prove that $\mathcal{M}_{s,t}$ and $\mathcal{L}_{s,t}$ have the same finite dimensional distributions for all fixed $s<t$. We will outline how we check that for a set $F=\{(x_{1},y_{1}),(x_{2},y_{2})\}$, the two-point distribution $\mathcal{M}_{0,1}|_{F}$ matches that of $\mathcal{L}$. The general case uses the same idea. For $s\in[0,1]$ define $\mathcal{M}^{s}=\mathcal{L}_{0,s}\diamond\mathcal{M}_{s,1},$ so that $\mathcal{M}^{0}=\mathcal{M}_{0,1}$ and $\mathcal{M}^{1}=\mathcal{L}_{0,1}$. We will aim to show that with respect to the $L^{1}$-Wasserstein metric $d_{W}$, for $s<t=s+r$,
@@
This type of compatibility can be viewed as roughly corresponding to the event where in a candidate landscape $\mathcal{M}^{t}$, if $\pi_{i},i=1,2$ are the geodesics from $(x_{i},0)$ and $(y_{i},1)$, then either $\pi_{1}(t),\pi_{2}(t)$ are far away from each other ($A_{1},A_{2}$ are separated), or else $\pi_{1}(t)=\pi_{2}(t)$ and the geodesics coalesce together well before time $s$ ($A_{1},A_{2}$ overlap and the functions $\mathcal{L}_{0,s}(x_{1},\cdot),\mathcal{L}_{0,s}(x_{2},\cdot)$ differ by a constant). In other words, the coalescence time for the geodesics $\pi_{1},\pi_{2}$ occurs far away from the interval $[s,t]$. If $\mathcal{M}^{s},\mathcal{M}^{t}$ were true directed landscapes, we could prove that this compatibility event holds with probability $1-r^{1-o(1)}$. For pre-landscapes, we will show the weaker bound $1-r^{2/3+\delta}$ for a fixed $\delta>0$, which still gives (11). In practice, we will also work with a stationary version of $\mathcal{M}^{s}|_{F},\mathcal{M}^{t}|_{F}$ in order to facilitate technical computations.
Remark 1.9.
The proof of Theorem 1.1 does not actually use much precise information about the directed landscape $\mathcal{L}$. Indeed, all we really need is that $\mathcal{L}$ satisfies the three conditions in Theorem 1.1 along with the following stationarity property, used to establish the $(1-r^{2/3+\delta})$-bound discussed above.
Property $\ast$:
@@
Future work.
One challenge in KPZ is constructing limit objects with boundaries and boundary conditions. In this direction, a major breakthrough was recently made by X. Zhang [Zha25], where the KPZ fixed point was constructed for functions on the half-line $0,\infty)$, with a source at the point 0. In forthcoming work, we build on the ideas of the present paper construct the directed landscape on the half-space $0,\infty)\times\mathbb{R}$ using Zhang’s half-space fixed point as the starting point. The idea is similar to the suggested construction of the full-space directed landscape in Remark 1.9. One key step is first constructing the half-space stationary horizon, which implies an analogue of Property $\ast$ in the half-space setting.
Acknowledgments. The authors thank Jeremy Quastel for several useful discussions regarding convergence of general AEPs. The authors also thank Amol Aggarwal for pointing out the argument in [ACH24b, Appendix D.3], and the references [BGRS02, BGRS19], and Ivan Corwin, Daniel Remenik, Milind Hegde, and Shalin Parekh for useful comments on an earlier version of the paper.
@@ 2.1 The directed landscape and the KPZ fixed point
Let $\operatorname{UC}$ be the set of all upper semicontinuous functions from $\mathbb{R}\to\bar{\mathbb{R}}$. A function $f$ is in $\operatorname{UC}$ if and only if its hypograph
@@
in the Hausdorff metric on closed subsets of $\mathbb{R}\times[-1,1]$, where $E:\mathbb{R}\times\bar{\mathbb{R}}\to\mathbb{R}\times[-1,1]$ is the map $E(x,s)=(x,e^{-|x|}s/(1+|s|)),$ and the $d_{H}$ metric on the right-hand side above is the usual Hausdorff metric. It is straightforward to check that $\operatorname{UC}$ is a compact Polish space. Setting some notations, we write $\delta_{x}\in\operatorname{UC}$ for the narrow wedge at $x$, i.e., the function with $\delta_{x}(x)=0$ and $\delta_{x}(y)=$-$\infty$ for $y\neq x$. For $f\in\operatorname{UC}$ and $K\subset\mathbb{R}$ we also write $f|_{K}\in\operatorname{UC}$ for the function equal to $f$ on $K$, and -$\infty$ on $K^{c}$.
@@
for all functions $f,-g\in\operatorname{UC}$ with at most linear growth at $\pm\infty$ through a Fredholm determinant formula. This characterizes the law of $\mathfrak{h}_{t}f$ for fixed $t,f$, which then characterizes the law of $(\mathfrak{h}_{t}f)_{t\geq 0}$ through the Markov property. The formula for the transition probabilities (14) is tractable enough to prove that for any $f$ with at most linear growth at $\pm\infty$, the map $(x,t)\mapsto\mathfrak{h}_{t}f(x)$ is Hölder-$1/2$- in $x$ and Hölder-$1/3$- in $t$.
@@
As already alluded to, we can alternately describe the KPZ fixed point directly in terms of the directed landscape. Indeed, for $f\in\operatorname{UC}$, define $\mathfrak{h}_{0}f=f$, and $\mathfrak{h}_{t}f$ for $t>0$ set $\mathfrak{h}_{t}f:=f\diamond\mathcal{L}_{0,t}$ (i.e.  equation (6)). Then $(\mathfrak{h}_{t}f)_{t\geq 0}$ is a time-homogeneous Markov process on $\operatorname{UC}$ whose transition probabilities match the Fredholm determinant formula for (14), at least for initial conditions with linear growth. Moving forward we consider (6) as defining the law of the KPZ fixed point $(\mathfrak{h}_{t}f)_{t\geq 0}$, which allows us to accommodate $f$ which may blow up faster than linearly at $\pm\infty$. Equation (6) also defines one particular coupling of the KPZ fixed point simultaneously from all different initial conditions. We refer to this as the landscape coupling.
@@ 2.2 Basic properties of the directed landscape
We say that $d$ is a directed metric of negative sign if d-d is a directed metric of positive sign. Many natural KPZ models are directed metrics but not metrics, and moreover, unlike the class of metrics, the class of directed metrics is preserved under the kinds of rescaling typical seen in the KPZ class. If we extend the directed landscape to equal -$\infty$ on $\mathbb{R}^{4}\setminus\mathbb{R}^{4}_{\uparrow}$ then it becomes a directed metric (of negative sign) on $\mathbb{R}^{2}$.
@@ 2.3 The stationary horizon
Here $f[i\to x]$ is a last passage value from $(i,-\infty)$ to $(1,x)$ in the sense of [DV24], and $|\pi|_{f}$ can be thought of as the length of a path associated to $\pi$. Now let $a_{1}<\dots<a_{k}$ and let $\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{k}$ be a family of independent two-sided Brownian motions where each $\mathcal{B}_{i}$ has slope $a_{i}$. Define $\mathcal{R}=\{\mathcal{R}_{i}\}_{i=1}^{k}$ by
@@ 3.1 Continuity and metric composition
on $\mathbb{Q}^{2}$. Using the continuity of $\mathcal{M}$, and Corollary 3.6 which controls the location of the argmax in the metric composition, we can deduce that both sides of (21) are continuous, and are also continuous in $s,r$, and $t$. Thus we can extend (21) to $\mathbb{R}^{2}$, and to hold for any $s<r<t$ simultaneously. ∎
@@ 3.2 The law of $\mathcal{M}_{0,1}$ and an exchange strategy
$(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{M}_{0,1}(y_{i})\}_{i=1}^{k})\stackrel{{\scriptstyle d}}{{=}}(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{L}_{0,1}(y_{i})\}_{i=1}^{k}).$ (22)
@@
By 2.5 and Corollary 3.6, with probability at least $1-C\exp(-cM)$, the bound in 2.5 holds for both $\mathcal{M},\mathcal{L}$ with some $R<M^{2/3}$. Also, with probability $1-C\exp(-cM)$, we have $|\mathcal{R}_{i}(x)-\mathcal{R}_{i}(0)-a_{i}x|<|x|+M$ for all $x\in\mathbb{R},i=1,\dots,k$. Let $\mathcal{E}$ denote the intersection of the above two events. Then on $\mathcal{E}$, we have that
@@ 3.3 The coupling
We now define the coupling to be used in 3.9. Below we fix $s<t=s+r\in[\theta,1-\theta]$, and throughout assume r=tsr=t-s is sufficiently small. We start with independent copies of $\mathcal{R}$, $\mathcal{L}_{0,s}$, and $\mathcal{M}_{t,1}$, and consider the following subsets of $\mathbb{R}$. For each $i\in\llbracket 1,k\rrbracket$, denote $H_{i}=\mathcal{R}_{i}\diamond\mathcal{L}_{0,s},$ and
@@
We let $A_{i}$ be the $\log^{10}(r)r^{2/3}$-neighborhood of $\tilde{A}_{i}$. For each $\in\llbracket 1,k\rrbracket, iji\neq j$, we let $\tilde{D}_{i,j}$ be the set where $H_{i}-H_{j}$ is non-constant, i.e.,
[…]
@@
There is a constant $\delta>0$, such that for all sufficiently small $r>0$ and $\in\llbracket 1,k\rrbracket, iji\neq j$, we have $\mathbb{P}[A_{i}\cap A_{j}\cap D_{i,j}=\emptyset]>1-r^{2/3+\delta}$.
@@
Let $\mathcal{E}_{*}$ be the event where $A_{i}\cap A_{j}\cap D_{i,j}=\emptyset$, for each $\in\llbracket 1,k\rrbracket, iji\neq j$. Note that for each $i\in\llbracket 1,k\rrbracket$, (using the shape bounds in 2.5, Corollary 3.6 on $\mathcal{L},\mathcal{M}$) $A_{i}$ consists of finitely many open intervals $A_{i,1},\dots,A_{i,\ell(i)}$, each of length at least $2\log^{10}(r)r^{2/3}$.
@@
On $\mathcal{E}_{*}^{c}$, set $F=0$ on $\bigcup_{i=1}^{k}A_{i}$ and -$\infty$ off this set. Next, for any interval $A_{i,j}$ and a function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ , we say that $f$ is suitable on $A_{i,j}$, if $f-H_{i}$ is constant on $I$. Sort the intervals $A_{i,j},i=1,\dots,k,j=1,\dots,\ell(i)$ into a list $I_{1},\ldots,I_{n}$ so that the $I_{j}$ are ordered by their left endpoint (ordering arbitrarily if left endpoints agree). We define $F$ on $I_{1},\dots,I_{n}$ inductively.
@@
It remains to check the last condition. By Theorem 2.9 the law of each $H_{i}$ is a Brownian motion with drift $a_{i}$. Then for any $M>1$, with probability $1-C\exp(-cM)$ we have that $|H_{i}(x)-H_{i}(y)|<M\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for each $i\in\llbracket 1,k\rrbracket$ and $x,y\in[-M,M]$. Also, by 2.5/Corollary 3.6, with probability $1-C\exp(-cM)$, we have that $\bigcup_{i=1}^{k}A_{i}\subset[-M,M]$. Thus the last condition holds from the construction of $F$. The measurability is immediate from construction. ∎
@@ 3.4 Removing the stationary initial condition
As $\epsilon\to 0$, the second bracketed term above remains bounded even after we take a supremum over $x$, since $\mathcal{R}_{i}$ is a Brownian motion with drift $i\leq k$; while by straightforward algebra the first term goes to -$\infty$ uniformly over $x$ with $|x-x_{i}|>\epsilon$. This yields (29). For the bound (30), observe that for any $f:\mathbb{R}\to\mathbb{R}^{n}$ and $x<x_{i}<y$ we have
@@
Using these inequalities with $f=\mathcal{B}+\phi_{\epsilon}$ yields (30) since $\mathcal{B}$ is continuous and each $\phi_{\epsilon,j}$ is non-decreasing on $x_{i}-\epsilon,x_{i}]$ and non-increasing on $x_{i},x_{i}+\epsilon]$. ∎
@@
We will show that (for some $\delta>0$) these two probabilities are $C\log^{44}(r)r^{4/3}$ and $r^{\delta}$, respectively.
@@
$\tilde{\mathcal{B}}_{2}(x)=\mathcal{B}_{1}(x)+\sup_{z\leq x}\mathcal{B}_{2}(z)-\mathcal{B}_{1}(z).$
@@
On the other hand, again using Lemma 2.8 and the stationarity of $x\mapsto\mathcal{A}_{2}(x)+x^{2}$, we can deduce that $\mathbb{P}[\mathcal{E}_{1}^{\prime}\cap\mathcal{E}_{2}^{\prime}]<Cr^{4/3}\log^{44}(r)$, by replacing $\mathcal{M}_{1}-\mathcal{M}_{1}(z_{+})$ on $z_{+}-1,z_{+}+1]$ with a Brownian motion $\mathcal{B}_{3}$ with $\mathcal{B}_{3}(z_{+})=0$ with drift in $C|\log(r)|,C|\log(r)|]$, and using Lemma 4.3 below (with $X=\mathcal{B}_{3}(z_{+}-\cdot),Y=\mathcal{B}_{1}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),Z=\mathcal{B}_{2}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),\epsilon=2\log^{10}(r)r^{1/3}$ and $D=C|\log(r)|$).
@@
Take any $D>1$ and $0<\epsilon<1$. Let $X,Y,Z$ be three independent Brownian motions, each with drift in [D,D][-D,D]. Then we have
@@
since $\tilde{\mathcal{B}}_{2}\geq\mathcal{B}_{2}$. Now, define the event $\mathcal{E}_{3}^{\prime}$ by
@@
We again use Lemma 2.8 and the stationarity of Airy2 plus a parabola, to replace $\mathcal{M}_{2}^{\prime}$ with a Brownian motion with drift in $C|\log(r)|,C|\log(r)|]$, conditioned on an event of probability at least $cr^{10}$. Then by Lemma 4.5 below (with $D=C|\log(r)|,Y=\mathcal{M}_{2}^{\prime}(\cdot+z_{+})-\mathcal{M}_{2}^{\prime}(z_{+}),X=\mathcal{B}_{2}(\cdot+z_{+})-\mathcal{B}_{2}(z_{+})$, and $\epsilon=2\log^{10}(r)r^{1/3}$) the above probability is at most $Cr^{\delta}$ for some absolute constant $\delta>0$. Putting everything together gives that
@@
There exists $\delta>0$ such that the following holds. Take any $D>1$ and $0<\epsilon<1$, with $D<\epsilon^{-1/4}$. Let $X$ and $Y$ be two independent Brownian motions, each with drift in [D,D][-D,D]. Let $\mathcal{E}$ be an event measurable with respect to $Y$, such that $\mathbb{P}[\mathcal{E}]>\epsilon^{100}$. Then
@@
For each $i\in\llbracket 1,2m\rrbracket$, we let $W^{\prime}_{i}$ be a Brownian motion with drift $2^{-(i+1)/2}D$, conditional on $W^{\prime}_{i}\leq 0$ in $[0,1]$. Fix $(i_{1},\cdots,i_{m})\in\Phi$. By the claim above, for each $j\in\llbracket 1,m\rrbracket$, $W^{\prime}_{i_{j}}$ is stochastically dominated by the process $x\mapsto 2^{(i_{j}+1)/2}X(2^{-i_{j}-1}x)$ conditional on $\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}^{\prime}$. Therefore
@@
for any II of length $r^{2/3}$, $I\subset[-|\log(r)|,|\log(r)|]$. Taking a union over at most $3r^{-2/3}\log(r)$ covering $|\log(r)|,|\log(r)|]$ and using Lemma 4.1 yields the result. ∎